The present value of a series of payments of $3 at the end of every eight years, forever, is
equal to $9.5. Calculate the effective annual rate of interest.
1
Expert's answer
2020-10-06T18:29:38-0400
"\nu=\dfrac{1}{1+i}""PV=3\cdot\nu^8 +3\cdot\nu^{16}+...=\dfrac{3\cdot \nu^8}{1-\nu^8 }=9.5""\nu^8=\dfrac{9.5}{12.5}=0.76""\dfrac{1}{1+i}=\sqrt[8]{0.76}""i=\dfrac{1}{\sqrt[8]{0.76}}-1\approx0.0349""i=3.49\%"